OpenStudy (anonymous):

write the nth term of the sequence 1, 4, 9, 16

OpenStudy (anonymous):

\[\sum_{k=1}^{n}2(k-1)+1\]

OpenStudy (anonymous):

What does that notation mean?

myininaya (myininaya):

the nth term is n^2

myininaya (myininaya):

since 1^2=1, 2^2=4, 3^2=9, 4^2=16,...,n^2

OpenStudy (anonymous):

so what would the nth term be?

OpenStudy (anonymous):

oh n^2 ok

OpenStudy (anonymous):

What about the sequence 3, 11, 31, 69, 131

OpenStudy (anonymous):

anyone there

OpenStudy (anonymous):

myininaya are you there?

myininaya (myininaya):

I'm trying to see the pattern

OpenStudy (anonymous):

The first differences would be 3, 20, 38, 62. The second differences would be 12, 18, 23 and 30. The third difference would be 6 for all of them

OpenStudy (anonymous):

Oh I meant 24 not 23

OpenStudy (anonymous):

Are you still there?

OpenStudy (anonymous):

That's funny. Both answers are correct. My answer was that the nth term was the sum of the first n odd natural numbers. So the first number is 1, the second is 1+3=4, the third is 1+3+5=9, the fourth is 1+3+5+7 = 16

OpenStudy (anonymous):

polpak did you get my second question?

OpenStudy (anonymous):

Yeah, looking at it, but these things are kinda lame cause there's a lot of different ways you can approach it.

OpenStudy (anonymous):

ok. tell me when you have an answer

OpenStudy (anonymous):

I don't get this stuff.